### v1

@@ Abstract
keywords:
symmetric $\alpha$-stable, stability problem, time-dependent coefficients, quantitative error estimates, non-Lipschitz coefficient, pathwise uniqueness
@@ 1 Introduction
To place our problem in context, it is helpful to first recall the classical case of Brownian motion $(\alpha=2)$. In this setting, the seminal work of Yamada and Watanabe [3] established the pathwise uniqueness of solutions under the condition that the diffusion coefficient $\sigma$ is $(1/2)$-Hölder continuous. The study of pathwise uniqueness has been actively pursued for SDEs driven by symmetric $\alpha$-stable processes as well. For the one-dimensional case, foundational results were established by Komatsu [4] and Bass et al. [5]. Their work demonstrated that pathwise uniqueness holds under a Hölder continuity condition on the coefficient $\sigma$ of order $1/\alpha$, which is the natural analogue of the Brownian case. This line of research was extended to the multi-dimensional setting by Tsuchiya [6], who also considered non-Lipschitz coefficients. Further generalizations have explored different aspects of the problem. For instance, Fournier [7] established uniqueness conditions under various assumptions on the process’s asymmetry and the coefficients’ regularity and monotonicity. The case where the stability index is small ($\alpha\in(0,1]$) was specifically addressed by Tsukada [8]. Beyond pathwise uniqueness, Kulik [9] investigated the existence and properties of the unique weak solution, including the Gaussian bounds of its density function. These studies on the existence and uniqueness of solutions naturally lead to the stability problem, which questions how the solution path changes when the coefficients of the SDE are perturbed. This problem is closely related to the analysis of numerical schemes, where, for instance, a series of studies on the convergence rate of Euler-Maruyama-type approximations have provided quantitative estimates under various coefficient regularities ([10], [11], [12], [13] and [14], et al.).
@@
While the zero-drift case is well-understood, the stability problem for non-zero drift is more complex. A notable recent contribution is by Hao and Wu [19], who addressed stability for a non-zero (distributional) drift, but restricted their analysis to the case of a constant jump coefficient ($\sigma=1$).
@@
$\mu_{x_{0},\sigma}^{\alpha,t}(dy)\coloneqq\frac{1}{t^{1/\alpha}\sigma^{1/\alpha}(x_{0})}g^{(\alpha)}\left(\frac{y-x_{0}}{t^{1/\alpha}\sigma^{1/\alpha}(x_{0})}\right)dy,$
@@
We denote this relationship by $g^{(\alpha)}(y)\asymp(|y|\vee 1)^{-\alpha}$. This two-sided estimate ensures that our measure $\mu_{x_{0},\sigma}^{\alpha,t}$ assigns less weight to regions far from the initial point $x_{0}$, naturally decreasing as the distance increases. The corresponding $L^{p}\left(\mu_{x_{0},\sigma}^{\alpha,t}\right)$-norm for the coefficient difference is then given by
$\|f\|_{L^{p}\left(\mu_{x_{0},\sigma}^{\alpha,t}\right)}\coloneqq\left(\int|f(y)|^{p}\mu_{x_{0},\sigma}^{\alpha,t}(dy)\right)^{1/p},$ (1.3)
@@
While these previous results have provided crucial insights, their applicability remains significantly limited by the prevailing assumptions of a zero drift term and time-homogeneous coefficients. Overcoming these restrictions is not merely a matter of technical generalization but a necessary step to align the theory with both practical applications and deeper mathematical challenges. Many real-world phenomena modeled by SDEs—from mean-reverting financial assets to complex systems in physics and biology—exhibit a deterministic tendency or "drift" that fundamentally governs their qualitative behavior. A robust stability theory must therefore be able to incorporate non-zero drift to be practically meaningful. Similarly, the assumption of time-constant coefficients is an oversimplification for most non-autonomous systems, where environmental factors like volatility in finance or external forces in engineering naturally vary over time. The inclusion of these features, however, presents substantial technical hurdles. The introduction of a drift term, in particular, creates additional integral terms within the classical Komatsu-type analysis that cannot be controlled by the estimates sufficient for the zero-drift case. Furthermore, handling time-dependent coefficients requires a conceptual shift in how the distance between functions is measured; a transient, large difference between coefficients should not necessarily imply instability, a subtlety that cannot be captured by simple supremum norms. This necessitates the development of a more flexible, integral-based measure of distance that is sensitive to the underlying dynamics of the process.
@@
Our work directly confronts these challenges by developing a unified framework that successfully incorporates both non-zero drift and time-dependent coefficients. This framework integrates the pioneering method of Komatsu [4] for pathwise uniqueness with transition density estimates from Knopova and Kulik [20], using a smooth approximation $u_{\delta,\varepsilon}$ of the function $u(x)=|x|^{\alpha-1}$. The key technical innovation of this paper is the extension of this method to handle SDEs with non-zero and time-dependent drift terms. This extension is non-trivial. While the original framework was effective for the zero-drift case (see [1]), the introduction of a drift term—especially a time-dependent one—creates additional terms that cannot be controlled by existing estimates. Our main contribution is twofold. First, we establish a new, sharper estimate for the derivative $|u_{\delta,\varepsilon}^{\prime}|$ (see Inequality (3.3)). This is the crucial technical tool required to bound the new terms arising from the drift. Second, we introduce a strategic optimization procedure for the auxiliary parameters $\varepsilon>0$ and $\delta>1$. These parameters are chosen not as constants, but as functions of the coefficient differences themselves. This dynamic choice is essential for balancing the diverging terms in our stability inequality, ultimately leading to the explicit convergence rates in our main theorem. This refined approach enables a unified treatment of both drift and jump coefficients, allowing us to derive quantitative stability bounds that hold beyond the zero-drift and time-independent setting. Furthermore, we apply this stability analysis to investigate convergence in probability, leveraging properties of quasi-martingales [21] to establish a rate of uniform convergence between solution paths.
@@ 2.2 The symmetric $\alpha$-stable process
The driving noise in our SDE is a symmetric $\alpha$-stable process $Z=\{Z_{t}\}_{t\in[0,T]}$ with a stability index $\alpha\in(1,2)$. We summarize its essential properties below. The jump of the process at time $t>0$ is denoted by $\Delta Z_{t}\coloneqq Z_{t}-Z_{t-}$, with $\Delta Z_{0}\coloneqq 0$. The process $Z$ can be characterized by its jump measure. The associated Poisson random measure $N$ on $\mathcal{B}([0,T])\times\mathcal{B}(\real\setminus\{0\})$ is defined by counting the jumps: $N(t,F)\coloneqq\sum_{0<s\leq t}1_{F}(\Delta Z_{s})$ for $F\in\mathcal{B}(\real\setminus\{0\})$. The intensity of these jumps is described by the Lévy measure $\nu$, which is given by the expectation $\nu(F)\coloneqq\mathbb{E}[N(1,F)]$. Throughout this paper, the Lévy measure of $Z$ is $\nu(dz)=c_{\alpha}|z|^{-1-\alpha}dz$. The compensated Poisson random measure, denoted by $\tilde{N}$, is then defined as the difference $\tilde{N}(\mathrm{d}z,\mathrm{d}s)\coloneqq N(\mathrm{d}z,\mathrm{d}s)-\nu(\mathrm{d}z)\mathrm{d}s$.
@@
The principal part $p_{t}^{0}(x_{0},\cdot)$ is the "frozen-coefficient" approximation, defined by
@@
As a direct consequence of this decomposition $p_{t}=p_{t}^{0}+r_{t}$ and the estimate on the residue term, it follows that the density $p_{t}$ is of the same order as its principal part $p_{t}^{0}$. Indeed, this is stated more directly as a two-sided estimate in Theorem 2.5 in [20] of, which shows that for some constants $m,M>0$,
@@ 3 Distance between solutions in the $L^{\alpha-1}(\Omega)$-norm
In this section, we present our first main result: a quantitative stability estimate for the solutions of SDE (1.1). Our result generalizes previous findings in two significant directions. First, whereas Hashimoto and Tsuchiya [18] measured the distance between coefficients using the supremum norm, our estimate is formulated in terms of the more refined $L^{p}\left(\mu_{x_{0},\sigma}^{\alpha,T}\right)$-norm introduced in the previous section. Second, our framework incorporates non-zero drift terms, extending the zero-drift analysis from [1]. The proof strategy involves a delicate balancing act. We will employ Komatsu’s method, which uses an auxiliary function $u_{\delta,\varepsilon}$ controlled by parameters $\varepsilon$ and $\delta$. Applying Itô’s formula yields a fundamental inequality, but this inequality contains coefficients that diverge as $\varepsilon\to+0$ or $\delta\to\infty$. The core idea of our proof is to strategically choose these parameters as functions of the norms measuring the differences between the coefficients. This optimization process balances the diverging terms against the smallness of the errors, and the explicit convergence rates presented in the following theorem are a direct consequence of this procedure.
We are now in a position to state our first main result. The following theorem provides an explicit bound on the $L^{\alpha-1}(\Omega)$ distance between the solution $X$ of our baseline SDE (1.1) and the solution $\tilde{X}$ of a more general SDE with time-dependent coefficients. As a consistency check, we note that in the special case where $b=\tilde{b}=0$, our result recovers the convergence rate established in [1, Theorem 31].
@@
    Jump coefficients: The coefficient $\sigma$ of $X$ is strictly positive, bounded, and Hölder continous:
@@
    The distances between the coefficients are small. These distances are measured by the quantities $B$ and $S$, defined by integrals weighted by the transition probability density $p_{t}(x_{0},\cdot)$ of the baseline process $X$ (as described in (2.1):
    $B\coloneqq\int_{0}^{T}\int\left|b(y)-\tilde{b}(s,y)\right|p_{s}\left(x_{0},y\right)\mathrm{d}s\mathrm{d}y<1,\quad S\coloneqq\left(\int_{0}^{T}\int\left|b(y)-\tilde{b}(s,y)\right|^{\alpha}p_{s}\left(x_{0},y\right)\mathrm{d}s\mathrm{d}y\right)^{1/\alpha}<1.$
@@
We add several remarks concerning the statement and assumptions of Theorem 3.1.
Remark 1.
@@
    (On the measure of coefficient distances) In our main estimate, the distances between the coefficients are measured by the quantities $B$ and $S$. This approach, which defines the distance as an integral weighted by the transition density $p_{s}(x_{0},\cdot)$, is a crucial feature of our framework for two reasons. First, it is specifically designed to handle time-dependent coefficients. A simpler measure, such as a supremum norm $\sup_{t\in[0,T]}\left\|b(\cdot)-\tilde{b}(t,\cdot)\right\|_{\infty}$, would require the coefficients to be uniformly close everywhere and at all times, which is a very restrictive condition. Our integral formulation allows for the pointwise differences to be large in regions where the process $X$ is unlikely to be (i.e., where $p_{t}(x_{0},\cdot)$ is small), as long as the overall weighted average remains small. This makes our stability result applicable to a much wider and more practical class of perturbations. Second, this weighted measure effectively localizes the error analysis. Since the transition density $p_{t}(x_{0},\cdot)$ is concentrated around the initial point $x_{0}$, especially for small times $t$, our measure focuses the analysis on the region most relevant to the process’s behavior. For the time-homogeneous case discussed in Corollary 5.1, these quantities $B$ and $S$ naturally correspond to the weighted $L^{p}$ norms, where the measure itself is constructed by integrating the transition density over time.

### v2

@@ Abstract
Keywords: symmetric $\alpha$-stable, stability problem, non-Lipschitz coefficient, pathwise uniqueness, error estimation
2020 Mathematics Subject Classification: 60H10; 60G52; 60H07
@@ 1 Introduction
To place our problem in context, it is helpful to first recall the classical case of Brownian motion $(\alpha=2)$. In this setting, the seminal work of Yamada and Watanabe [26] established pathwise uniqueness of solutions under the condition that the diffusion coefficient $\sigma$ is $(1/2)$-Hölder continuous. The study of pathwise uniqueness has been actively pursued for SDEs driven by symmetric $\alpha$-stable processes as well. For the one-dimensional case, foundational results were established by Komatsu [16] and Bass et al. [2]. Their work demonstrated that pathwise uniqueness holds under a Hölder continuity condition on the coefficient $\sigma$ of order $1/\alpha$, which is the natural analogue of the Brownian case. This line of research was extended to the multi-dimensional setting by Tsuchiya [23], who also considered non-Lipschitz coefficients. Further generalizations have explored different aspects of the problem. For instance, Fournier [5] established uniqueness conditions under various assumptions on the process’s asymmetry and the coefficients’ regularity and monotonicity. The case where the stability index is small ($\alpha\in(0,1]$) was specifically addressed by Tsukada [10]. Beyond pathwise uniqueness, Kulik [15] investigated the existence and properties of the unique weak solution, including the Gaussian bounds of its density function. These studies on the existence and uniqueness of solutions naturally lead to the stability problem, which questions how the solution path changes when the coefficients of the SDE are perturbed. This problem is closely related to the analysis of numerical schemes, where, for instance, a series of studies on the convergence rate of Euler-Maruyama-type approximations have provided quantitative estimates under various coefficient regularities ([7], [12], [19], [21] and [25], et al.).
@@
While the zero-drift case is well-understood, the stability problem for non-zero drift is more complex. A notable recent contribution is by Hao and Wu [6]. They established a quantitative stability estimate for SDEs with irregular drifts driven by a cylindrical $\alpha$-stable process, where the distance between solutions is measured in the total variation distance of the laws. Their framework accommodates a general class of multiplicative noise. Our work differs from and complements the results in two significant aspects. First, whereas their analysis focuses on perturbations in the drift coefficient, our framework treats perturbations in both the drift and the jump coefficients simultaneously. Second, and more fundamentally, we establish stability in terms of the pathwise $L^{\alpha-1}(\Omega)$ distance between two solutions, which is a stronger notion of convergence than the convergence of laws. Our result thus provides a different type of quantitative estimate that captures the path-by-path proximity of solutions, which is often crucial for applications such as numerical scheme analysis.
@@
$\mu_{x_{0},\sigma}^{\alpha,t}(\mathrm{d}y)\coloneqq\frac{1}{t^{1/\alpha}\sigma^{1/\alpha}(y)}g^{(\alpha)}\left(\frac{y-x_{0}}{t^{1/\alpha}\sigma^{1/\alpha}(y)}\right)\mathrm{d}y,$
@@
We denote this relationship by $g^{(\alpha)}(y)\asymp(|y|\vee 1)^{-1-\alpha}$. This measure $\mu_{x_{0},\sigma}^{\alpha,t}$ is constructed to be directly comparable to the principal part of the transition density, $p_{t}^{0}(x_{0},y)$ (see Subsetion 2.3). Indeed, both share the same fundamental structure, incorporating the space-dependent coefficient $\sigma$. A key feature of this measure is that it decays as the distance $|y-x_{0}|$ increases. This follows from the decay property of $g^{(\alpha)}$ and the assumption that $\sigma$ is uniformly bounded from above and below, which ensures the argument of $g^{(\alpha)}$ grows linearly with $|y-x_{0}|$. This decay property ensures that $L^{p}\left(\mu_{x_{0},\sigma}^{\alpha,t}\right)$-norm, defined below, effectively localizes the error analysis to the region near the initial point $x_{0}$.
$\|f\|_{L^{p}\left(\mu_{x_{0},\sigma}^{\alpha,t}\right)}\coloneqq\left(\int|f(y)|^{p}\mu_{x_{0},\sigma}^{\alpha,t}(\mathrm{d}y)\right)^{1/p},$ (1.3)
@@
While these previous results have provided crucial insights, their applicability remains significantly limited by the prevailing assumptions of a zero drift term and time-homogeneous coefficients. Overcoming these restrictions is not merely a matter of technical generalization but a necessary step to align the theory with both practical applications and deeper mathematical challenges. Many real-world phenomena modeled by SDEs—from mean-reverting financial assets to complex systems in physics and biology—exhibit a deterministic tendency or ”drift” that fundamentally governs their qualitative behavior. A robust stability theory must therefore be able to incorporate non-zero drift to be practically meaningful. Similarly, the assumption of time-constant coefficients is an oversimplification for most non-autonomous systems, where environmental factors like volatility in finance or external forces in engineering naturally vary over time. The inclusion of these features, however, presents substantial technical hurdles. The introduction of a drift term, in particular, creates additional integral terms within the classical Komatsu-type analysis that cannot be controlled by the estimates sufficient for the zero-drift case. Furthermore, handling time-dependent coefficients requires a conceptual shift in how the distance between functions is measured; a transient, large difference between coefficients should not necessarily imply instability, a subtlety that cannot be captured by simple supremum norms. This necessitates the development of a more flexible, integral-based measure of distance that is sensitive to the underlying dynamics of the process.
@@
Our work directly confronts these challenges by developing a unified framework that successfully incorporates both non-zero drift and time-dependent coefficients. This framework integrates the pioneering method of Komatsu [16] for pathwise uniqueness with transition density estimates from Knopova and Kulik [14], using a smooth approximation $u_{\delta,\varepsilon}$ of the function $u(x)=|x|^{\alpha-1}$. The key technical innovation of this paper is the extension of this method to handle SDEs with non-zero and time-dependent drift terms. This extension is non-trivial. While the original framework was effective for the zero-drift case (see [20]), the introduction of a drift term—especially a time-dependent one—creates additional terms that cannot be controlled by existing estimates. Our methodological approach extends the classical Komatsu-type argument to a more general setting. This extension is highly non-trivial, as the introduction of a drift term creates additional integral terms that cannot be controlled by existing estimates. The genuine novelty of this paper lies in two main contributions that overcome this challenge. First, and most critically, we establish a new, sharper derivative estimate for the mollified auxiliary function (inequality (3.3)). This inequality is the technical core of our work. It provides a refined bound that is essential for decomposing the drift term into a manageable, recursive part and a part related to the direct coefficient difference. Without this estimate, the application of Gronwall’s inequality to the recursive drift term would not be possible, and the entire stability argument would fail. Second, building on this new estimate, we introduce a strategic optimization procedure for the auxiliary parameters $\varepsilon>0$ and $\delta>1$. These parameters are chosen not as constants, but as functions of the coefficient differences themselves. This dynamic choice is essential for balancing the diverging terms in our stability inequality, ultimately leading to the explicit convergence rates in our main theorem. This refined approach enables a unified treatment of both drift and jump coefficients, allowing us to derive quantitative stability bounds that hold beyond the zero-drift and time-independent setting. Furthermore, we apply this stability analysis to investigate convergence in probability, leveraging properties of quasi-martingales [17] to establish a rate of uniform convergence between solution paths in probability. It is important to note the nature of our stability result. Our framework is intentionally asymmetric: we begin with a well-understood, time-homogeneous baseline process and quantify its stability under time-dependent perturbations. This approach models the common practical question of how a standard, established model behaves when subjected to more complex, non-autonomous influences. Consequently, our main theorem should be interpreted as a sensitivity estimate for a baseline model against a broad class of time-dependent perturbations.
@@ 2.2 The symmetric $\alpha$-stable process
The driving noise in our SDE is a symmetric $\alpha$-stable process $Z=\{Z_{t}\}_{t\in[0,T]}$ with a stability index $\alpha\in(1,2)$. We summarize its essential properties below. The jump of the process at time $t>0$ is denoted by $\Delta Z_{t}\coloneqq Z_{t}-Z_{t-}$, with $\Delta Z_{0}\coloneqq 0$. The process $Z$ can be characterized by its jump measure. The associated Poisson random measure $N$ on $\mathcal{B}([0,T])\times\mathcal{B}(\real\setminus\{0\})$ is defined by counting the jumps: $N(t,F)\coloneqq\sum_{0<s\leq t}1_{F}(\Delta Z_{s})$ for $F\in\mathcal{B}(\real\setminus\{0\})$. The intensity of these jumps is described by the Lévy measure $\nu$, which is given by the expectation $\nu(F)\coloneqq\mathbb{E}[N(1,F)]$. Throughout this paper, the Lévy measure of $Z$ is $\nu(\mathrm{d}z)=c_{\alpha}|z|^{-1-\alpha}\mathrm{d}z$. The compensated Poisson random measure, denoted by $\tilde{N}$, is then defined as the difference $\tilde{N}(\mathrm{d}z,\mathrm{d}s)\coloneqq N(\mathrm{d}z,\mathrm{d}s)-\nu(\mathrm{d}z)\mathrm{d}s$.
@@
The principal part $p_{t}^{0}(x_{0},\cdot)$ is the ”frozen-coefficient” approximation, defined by
@@
As a direct consequence of this decomposition $p_{t}=p_{t}^{0}+r_{t}$ and the estimate on the residue term, it follows that the density $p_{t}$ is of the same order as its principal part $p_{t}^{0}$. Indeed, this is stated more directly as a two-sided estimate in Theorem 2.5 in [14] of which shows that for some constants $m,M>0$,
@@ 3 Distance between solutions in the $L^{\alpha-1}(\Omega)$-norm
In this section, we present our first main result: a quantitative stability estimate for the solutions of SDE (1.1). Our result generalizes previous findings in two significant directions. First, whereas Hashimoto and Tsuchiya [8] measured the distance between coefficients using the supremum norm, our estimate is formulated in terms of the more refined $L^{p}\left(\mu_{x_{0},\sigma}^{\alpha,T}\right)$-norm introduced in the previous section. Second, our framework incorporates non-zero drift terms, extending the zero-drift analysis from [20]. The proof strategy involves a delicate balancing act. We will employ Komatsu’s method, which uses an auxiliary function $u_{\delta,\varepsilon}$ controlled by parameters $\varepsilon$ and $\delta$. Applying Itô’s formula yields a fundamental inequality, but this inequality contains coefficients that diverge as $\varepsilon\to+0$ or $\delta\to\infty$. The core idea of our proof is to strategically choose these parameters as functions of the norms measuring the differences between the coefficients. This optimization process balances the diverging terms against the smallness of the errors, and the explicit convergence rates presented in the following theorem are a direct consequence of this procedure. In addition, it is important to emphasize that the restriction $\alpha>1$ is not merely a consequence of the cited transition density theory introduced in Subsection 2.3, but is intrinsic to our Komatsu-type approach. Our proof is built upon an auxiliary function, $u_{\delta,\varepsilon}$, which is a mollified version of the function $u(x)=|x|^{\alpha-1}$. The entire framework relies on the exponent $\alpha-1$ being positive, which ensures that this underlying function is regular at the origin and suitable for measuring pathwise distance. The case $\alpha\leq 1$, known as the supercritical case, would require a fundamentally different methodology and is thus outside the scope of this paper.
We are now in a position to state our first main result. The following theorem provides an explicit bound on the $L^{\alpha-1}(\Omega)$ distance between the solution $X$ of our baseline SDE (1.1) and the solution $\tilde{X}$ of a more general SDE with time-dependent coefficients. As a consistency check, we note that in the special case where $b=\tilde{b}=0$, our result recovers the convergence rate established in [20, Theorem 31].
@@
    Jump coefficients: The coefficient $\sigma$ of $X$ is strictly positive, bounded, and Hölder continuous:
@@
    The distances between the coefficients are small. These distances are measured by the quantities $B$ and $S$, defined by integrals weighted by the transition probability density $p_{t}(x_{0},\cdot)$ of the baseline process $X$ (as described in (2.1)):
    $B\coloneqq\int_{0}^{T}\int\left|b(y)-\tilde{b}(s,y)\right|p_{s}\left(x_{0},y\right)\mathrm{d}s\mathrm{d}y<1,\quad S\coloneqq\left(\int_{0}^{T}\int\left|\sigma(y)-\tilde{\sigma}(s,y)\right|^{\alpha}p_{s}\left(x_{0},y\right)\mathrm{d}s\mathrm{d}y\right)^{1/\alpha}<1.$
@@
It is important to interpret the stability estimate of Theorem 3.1 as a measure of overall stability across the entire time horizon. The resulting estimate does not describe how the error evolves from moment to moment, but rather captures the total accumulated effect of the coefficient perturbations over the whole interval $[0,T]$. We add several remarks concerning the statement and assumptions of Theorem 3.1.
Remark 3.2.
@@
[…]
@@
  4. 4.
    (Asymmetric structure around a baseline model) The stability estimate is asymmetric. The roles of the baseline process $X$ and the perturbed process $\tilde{X}$ are not interchangeable. The distances between coefficients, $B$ and $S$, are always measured using the transition probability density $p_{t}(x_{0},\cdot)$ of the baseline process $X$. This structure is intentional and reflects the core purpose of the theorem: to provide a sensitivity analysis of a well-understood, time-homogeneous model ($X$) against a wide range of time-dependent perturbations ($\tilde{X}$). This framework is particularly relevant in applications where one wants to assess the robustness of a standard model against realistic, non-autonomous environmental changes.
  5. 5.
    (On the solution framework) Our main theorem provides a stability estimate between two solutions, $X$ and $\tilde{X}$. We clarify here the precise solution concepts and the underlying assumptions on their existence.
    • For the baseline process $X$: Under our assumptions, the existence of a weak solution and its transition probability density is guaranteed by the work of Knopova and Kulik [14]. We do not assume pathwise uniqueness for $X$, as it is known to fail if the Hölder exponent $\eta<1/\alpha$ (see Bass et al. [2]).
    • For the perturbed process $\tilde{X}$: Our theorem is formulated under the hypothesis that a strong solution $\tilde{X}$ exists. The question of strong well-posedness for SDEs driven by stable processes with Hölder continuous coefficients is a highly non-trivial problem at the forefront of current research. To place this in context, recent significant results have been established under different sets of assumptions. For instance, Chen, et al. [3] proved strong well-posedness when the diffusion coefficient $\tilde{\sigma}$ is Lipschitz continuous and the drift $b$ is Hölder continuous (with exponent $\beta>1-\alpha/2$). In the Brownian motion case $(\alpha=2)$, Hutzenthaler and Jentzen [9] have shown the existence of strong solutions even when the diffusion coefficient is merely Hölder continuous. However, extending such results to the $\alpha$-stable case with a Hölder continuous diffusion coefficient remains a challenging open problem.
    Therefore, our main result should be interpreted as follows: if a strong solution $\tilde{X}$ to the perturbed SDE exists on a given probability space, then for any weak solution $X$ of the baseline SDE on the same space, their distance is bounded by the explicit rate given in Theorem 3.1. Proving the existence of $\tilde{X}$ under our specific, time-dependent Hölder assumptions is a substantial research problem in its own right and is considered beyond the scope of this paper.